### v2

@@ 1. Introduction
The paper is organized as follows. In Section 2, we recall the basic notions that will appear throughout the paper. In Section 3, we explore the connection between exponential and $h$-dichotomies. Afterwards, in Section 4 we introduce the concepts of uniform $h$-noncriticality and $h$-expansiveness for an arbitrary invertible evolution family $T(t,s)$ and note that these can also be characterized in terms of uniform noncriticality and exponential expansiveness of $T_{h}(t,s)$. Finally, in Section 5 we obtain the main results of this paper. Firstly, we obtain the version of [15, Theorem 1] for invertible evolution families on Banach spaces (see Theorem 1) providing characterizations of exponential dichotomies. Using the other results of our paper, we then obtain the version of Theorem 1 for $h$-dichotomies as a simple corollary (see Theorem 2).
@@ 2. Preliminaries
In the particular case where $h(t)=e^{t}$, the notion of an $h$-bounded growth coincides with the classical notion of a bounded growth (see [12, p.334]). In addition, in this case $h$-bounded decay will be called bounded decay.
@@
In the particular case where $h(t)=e^{t}$, the notion of a $h$-dichotomy coincides with the notion of an exponential dichotomy as introduced by Henry (see [8, Definition 7.6.1]).
@@ 4. Noncritical uniformity and expansiveness
In the case $h(t)=e^{t}$, we will say that $\mathcal{T}$ is exponentially expansive. For evolution families arising from nonautonomous ordinary differential equations, this definition coincides with [15, Definition 5].
@@
In the case $h(t)=e^{t}$, we will say that $\mathcal{T}$ is uniformly noncritical. For evolution families arising from nonautonomous ordinary differential equations, this definition coincides with [15, Definition 4]. This notion was introduced for nonlinear systems by Krasovski [9] and later adapted to linear systems by Massera and Schäffer [11].
@@
$\begin{split}\|v\|&\leq\theta\sup\{\|T(s,h^{-1}(e^{t}))v\|:\ |\ln h(s)-\ln h(h^{-1}(e^{t}))|\leq C\}\\ &=\theta\sup\{\|T(h^{-1}(e^{u}),h^{-1}(e^{t}))\|:\ |\ln h(h^{-1}(e^{u}))-\ln h(h^{-1}(e^{t}))|\leq C\}\\ &=\theta\sup\{\|T_{h}(u,t)v\|:\ |u-t|\leq C\}.\end{split}$
We conclude that $\mathcal{T}^{h}$ is uniformly noncritical on $\ln h(a_{0}^{*}),\infty)$. The converse can be established similarly. ∎
@@ 5.1. Characterization of exponential dichotomies
The following is the first main result of our paper. Its proof is inspired by the proof of [2, Proposition 1, p.14].
@@
Here $T(s,a_{0}^{*})Z$ denotes the image of the subspace $Z$ under the action of $T(s,a_{0}^{*})$. Fix $s\geq a_{0}^{*}$ and $v\in\mathcal{S}(s)\setminus\{0\}$. Then
@@
and, consequently, $\varrho\leq Ke^{\mu C}\|v\|$ (where we used (2.1)). We conclude that
@@
where $D\geq 1$ is independent of $s$ and $v$. As in the proof of [2, Proposition 1, p.15], (5.5) (together with (5.3)) implies that
@@
where $B:=\theta^{-1}D$ and $\alpha:=-C^{-1}\ln\theta$. Indeed, take $v\in X$, $t\geq s\geq a_{0}^{*}$ and let $n\in\mathbb{N}_{0}$ be such that $s+nC\leq t<s+(n+1)C$. By (5.3),
@@
It follows from (2.1), (5.6) and (5.11) (together with [12, Lemma 4.2]) that $\mathcal{T}$ admits an exponential dichotomy on $a_{0}^{*},\infty)$ with respect to projections $P(t)\colon X\to\mathcal{S}(t)$ associated to (5.4). ∎
@@
    where $A\colon(a_{0},\infty)\to\mathcal{B}(X)$ is continuous, we can remove the assumption that $T(t,s)$ exhibits a bounded decay (see the argument in [2, p.13]).
@@
$\begin{split}\operatorname{Im}P(t)&=\left\{(f,Mf)\in X:\ \sup_{r\geq t}\|\bar{T}(r,t)(f,Mf)\|<+\infty\right\}\\ &=\left\{(f,Mf)\in X:\ \sup_{r\geq t}\|T(r-t)(f,Mf)\|<+\infty\right\}\\ &=\left\{(f,Mf)\in X:\sup_{r\geq 0}\|T(r)(f,Mf)\|=0\right\},\end{split}$

### v3

@@ 1. Introduction
The paper is organized as follows. In Section 2, we recall the basic notions that will appear throughout the paper. In Section 3, we explore the connection between exponential and $h$-dichotomies. Afterwards, in Section 4 we introduce the concepts of uniform $h$-noncriticality and $h$-expansiveness for an arbitrary invertible evolution family $T(t,s)$ and note that these can also be characterized in terms of uniform noncriticality and exponential expansiveness of $T_{h}(t,s)$. Finally, in Section 5 we obtain the main results of this paper. Firstly, we obtain the version of [16, Theorem 1] for invertible evolution families on Banach spaces (see Theorem 1) providing characterizations of exponential dichotomies. Using the other results of our paper, we then obtain the version of Theorem 1 for $h$-dichotomies as a simple corollary (see Theorem 2).
@@ 2. Preliminaries
In the particular case where $h(t)=e^{t}$, the notion of an $h$-bounded growth coincides with the classical notion of a bounded growth (see [20, Definition 2.1.(iv)]). We note that this requirement is equivalent to that in [13, p.334] which requires that there are $K\geq 0$ and $\mu\in\mathbb{R}$ such that (2.1) holds, as for $\mu\leq 0$ we have $e^{\mu(t-s)}\leq 1$ for $t\geq s$.
In addition, in the case where $(t)=e^{t}, hh$-bounded decay will be called bounded decay.
@@
In the particular case where $h(t)=e^{t}$, the notion of a $h$-dichotomy coincides with the notion of an exponential dichotomy as introduced by Henry (see [9, Definition 7.6.1]).
@@ 4. Noncritical uniformity and expansiveness
In the case $h(t)=e^{t}$, we will say that $\mathcal{T}$ is exponentially expansive. For evolution families arising from nonautonomous ordinary differential equations, this definition coincides with [16, Definition 5].
@@
In the case $h(t)=e^{t}$, we will say that $\mathcal{T}$ is uniformly noncritical. For evolution families arising from nonautonomous ordinary differential equations, this definition coincides with [16, Definition 4]. This notion was introduced for nonlinear systems by Krasovski [10] and later adapted to linear systems by Massera and Schäffer [12].
@@
$\begin{split}\|v\|&\leq\theta\sup\{\|T(s,h^{-1}(e^{t}))v\|:\ |\ln h(s)-\ln h(h^{-1}(e^{t}))|\leq C\}\\ &=\theta\sup\{\|T(h^{-1}(e^{u}),h^{-1}(e^{t}))\|:\ |\ln h(h^{-1}(e^{u}))-\ln h(h^{-1}(e^{t}))|\leq C\}\\ &=\theta\sup\{\|T_{h}(u,t)v\|:\ |u-t|\leq C\},\end{split}$
where in the second step, we made the change of variables $u=\ln h(s)$. We conclude that $\mathcal{T}^{h}$ is uniformly noncritical on $\ln h(a_{0}^{*}),\infty)$. The converse can be established similarly. ∎
@@ 5.1. Characterization of exponential dichotomies
The following is the first main result of our paper. Its proof is inspired by the proof of [3, Proposition 1, p.14].
@@
Here $T(s,a_{0}^{*})Z$ denotes the image of the subspace $Z$ under the action of $T(s,a_{0}^{*})$. Note that $\mathcal{S}(a_{0}^{*})=\mathcal{S}$ and $Z(a_{0}^{*})=Z$. Fix $s\geq a_{0}^{*}$ and $v\in\mathcal{S}(s)\setminus\{0\}$. Then
@@
and, consequently, $\varrho\leq Ke^{\mu C}\|v\|$ (where we used (2.1) with $h(t)=e^{t}$). We conclude that
@@
where $D\geq 1$ is independent of $s$ and $v$. As in the proof of [3, Proposition 1, p.15], (5.5) (together with (5.3)) implies that
@@
where $B:=\theta^{-1}D$ and $\alpha:=-C^{-1}\ln\theta$. Indeed, take $v\in\mathcal{S}(s)$, $t\geq s\geq a_{0}^{*}$ and let $n\in\mathbb{N}_{0}$ be such that $s+nC\leq t<s+(n+1)C$. By (5.3),
@@
It follows from (2.1), (5.6) and (5.11) (together with [13, Lemma 4.2]) that $\mathcal{T}$ admits an exponential dichotomy on $a_{0}^{*},\infty)$ with respect to projections $P(t)\colon X\to\mathcal{S}(t)$ associated to (5.4). ∎
@@
    where $A\colon(a_{0},\infty)\to\mathcal{B}(X)$ is continuous, we can remove the assumption that $T(t,s)$ exhibits a bounded decay (see the argument in [3, p.13]).
@@
$\begin{split}\operatorname{Im}P(t)&=\left\{(f,Mf)\in X:\ \sup_{r\geq t}\|\bar{T}(r,t)(f,Mf)\|<+\infty\right\}\\ &=\left\{(f,Mf)\in X:\ \sup_{r\geq t}\|T(r-t)(f,Mf)\|<+\infty\right\}\\ &=\left\{(f,Mf)\in X:\sup_{r\geq 0}\|T(r)(f,Mf)\|<+\infty\right\},\end{split}$